# Cost Functions in Economic Complexity

### v1

@@ Abstract
Economic complexity algorithms, including the Economic Complexity Index (ECI) and Non-Homogeneous Economic Fitness and Complexity (NHEFC), have proven effective at capturing the intricate dynamics of economic systems. Here, we present a fundamental reinterpretation of these algorithms by reformulating them as optimization problems that minimize specific cost functions. We show that ECI computation is equivalent to finding eigenvectors of the network’s transition matrix by minimizing the quadratic form associated with the network’s Laplacian. For NHEFC, we derive a novel cost function that exploits the algorithm’s intrinsic logarithmic structure and clarifies the role of its regularization parameter. Additionally, we establish the uniqueness of the NHEFC solution, providing theoretical foundations for its application. This optimization-based reformulation bridges economic complexity and established frameworks in spectral theory, network science, and optimization. The theoretical insights translate into practical computational advantages: we introduce a conservative, gradient-based update rule that substantially accelerates algorithmic convergence. Beyond advancing our theoretical understanding of economic complexity indicators, this work opens new pathways for algorithmic improvements and extends applicability to general network structures beyond traditional bipartite economic networks.
@@ 1 Introduction
Economic complexity measures provide valuable insights into the productive capabilities of countries and the structure of international trade [1, 2]. Two of the most widely used measures, the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC), have been instrumental in quantifying the economic diversification of nations and the ubiquity of products [3]. The foundational works by Hidalgo and Hausmann [4, 5] introduced ECI as a method to infer a country’s economic capabilities by analyzing the structure of the country-product network. Subsequently, Tacchella et al. [6, 7] proposed an alternative approach with the Economic Fitness and Complexity method, introducing a non-linear iterative framework to better capture the interplay between country diversification and product competitiveness [8]. Later refinements of EFC, such as the regularized Non-Homogeneous Economic Fitness and Complexity (NHEFC) method [9], aimed to ensure robust convergence and improve its applicability to empirical data.
@@
While these algorithms have demonstrated strong empirical relevance in understanding economic growth and trade networks, their mathematical foundations remain an active area of research. Previous studies [10, 11] have explored the connections between economic complexity indices and spectral methods in network science. For example, in [12] the authors show how ECI and EFC can be integrated into a unified measure incorporating information from both approaches, highlighting their structural similarities despite differences in algorithmic form. In [13, 14], the authors investigate the relation between economic complexity indices and standard notions of centrality, showing that these algorithms can provide valuable insight into the spectral properties of the underlying graph. More recently, ECI and EFC have also been shown to be formally equivalent to network centrality measures [15]. Extending these algorithms to arbitrary network structures, beyond the canonical bipartite formulation, has significantly broadened their applicability and interpretability. Therefore, a deeper theoretical understanding of these approaches is crucial to expanding their potential applications, as already demonstrated in fields such as ecology [16, 17, 18].
@@
Building upon these developments, recent studies have proposed a connection between the EFC algorithm and optimal transport theory [19], particularly in relation to the Sinkhorn-Knopp algorithm [20], which is used to compute doubly stochastic matrices. In this context, economic complexity indices can be interpreted as solutions to optimization problems analogous to those found in optimal transport [21, 22], where cost functions play a central role in defining the equilibrium distributions [23]. Notably, the iterative structure of the NHEFC algorithm resembles the scaling procedures used in transport problems, and its associated cost function exhibits a formal analogy with Kantorovich potentials [24], which characterize the optimal allocation of resources over networks. More generally, the interpretation of economic complexity algorithms as optimization problems [25, 26, 27] provides a unified framework that clarifies their mathematical structure, reveals properties of their solutions, and facilitates connections with other domains.
@@
This study demonstrates that both ECI and NHEFC can be reformulated as optimization problems, each minimizing a distinct cost function. Specifically, we show that the ECI algorithm minimizes an energy function derived from the network’s Laplacian matrix. In contrast, we prove that the cost function underlying NHEFC is strictly convex, thus providing a rigorous foundation for its optimization-based formulation. We also establish convergence to the fixed point—corresponding to the minimum of the cost function—in the case of regular graphs, and prove local convergence for general undirected and unweighted graphs. In addition, we propose a modification to the iterative algorithm that directly follows the gradient descent of the potential, significantly accelerating numerical convergence.
@@ 2.2 Previous work
In our previous work [15, 23], we described two complementary methods for quantifying economic complexity: the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC). Both approaches are based on the country–product binary matrix $\mathbf{M}$. In this matrix, the entry $M_{cp}$ is set to 1 if country $c$ significantly exports product $p$, and 0 otherwise, as determined by the Revealed Comparative Advantage (RCA) criterion [30].
@@
In contrast, EFC employs a non-linear map [6]. To ensure robust convergence, we used the regularized Non-Homogeneous EFC (NHEFC) given by [9]
@@
By introducing the adjacency matrix $\mathbf{A}$ and the diagonal matrix $\mathbf{K}$ with $k_{c}$ and $k_{p}$ on its diagonal, we can provide a mono-partite representation that recasts both ECI and NHEFC as iterative maps over a unified vector $\vec{V}$ of dimension $N=N_{c}+N_{p}$.
@@
thereby generalizing the approach, originally designed for bipartite structures, to any simple-graph structure. In Eq. (3), the notation $(\vec{V})^{-1}$ stands for taking the reciprocal of each element of the vector $\vec{V}$. This mono-partite formulation is a key generalization, allowing the ECI and NHEFC concepts to be applied to any arbitrary network structure, not just the bipartite country-product networks for which they were originally designed.
@@ 2.3 Cost Functions
In this section, we show that both ECI and NHEFC can be understood as optimization algorithms, as they can be reformulated as problems of minimizing specific cost functions. In many optimization settings, the cost (or objective) function is defined from the outset, and iterative procedures are devised to find its minimum. In our case, the process is reversed: starting from the iterative maps in Eq. (2) and Eq. (3), we reconstruct the corresponding cost functions by integration. The cost function associated with ECI relates directly to the standard eigenvector problem and can be interpreted within that framework. For NHEFC, the connection to optimization has been previously noted due to structural similarities between the fitness algorithm and the Sinkhorn-Knopp algorithm from optimal transport theory [31]. However, while prior work has identified this formal analogy [19], our study provides the first explicit derivation of NHEFC’s cost function. This explicit formulation enables us to prove the function’s convexity rigorously and subsequently develop a more efficient gradient-based algorithm. In the following, we derive the explicit expression of the cost function associated with the NHEFC algorithm. Full details of the derivations for both ECI and NHEFC are provided in Appendix A.
@@ 3 Results
In this section, we analyze the convergence properties of ECI and NHEFC. For ECI, properties such as the convexity of the potential and the existence of a fixed point follow directly from the fact that ECI is equivalent to computing the second eigenvector of the matrix $\mathbf{K}^{-1}\mathbf{A}$. The following part focuses on the convergence of NHEFC across different network structures, starting from a simple two-node undirected graph and extending to regular graphs and arbitrary topologies. The analysis establishes local convergence criteria near fixed points and explores the influence of the non-homogeneous parameter $\delta$ on the algorithm’s behavior. Finally, we introduce an alternative formulation of the fitness algorithm that follows the gradient of the potential function. This conservative formulation defines the most direct path to the fixed point and significantly reduces the number of iterations required for convergence.
@@ 3.1 Convergence properties of ECI
Since this fixed point (the constant vector) does not provide a meaningful solution of the problem of ranking countries or products, the original formulation of ECI [4] prescribes stopping the iteration after a fixed number of steps. As discussed in several works [12, 23, 1], the intermediate steps of the iteration retain information about the subdominant eigenvectors, particularly the second one. As a result, ECI can be obtained in two equivalent ways: (i) by iterating the map Eq. (2) for a fixed number of steps, or (ii) by directly computing the second eigenvector of the transition matrix $\mathbf{K}^{-1}\mathbf{A}$. This eigenvector maximizes the partitioning of the network, in line with the interpretation of ECI as a spectral algorithm for community detection [12].
3.2 Convergence properties of NHEFC
@@ 3.2.1 Uniqueness of solution via convexity analysis
Figure 1: Energetic landscape of the fitness algorithm. a) Cost function from Eq. (7) for regular complete graphs with varying number of nodes $N$ and $K=N-1$. The potential $U(V)$ is plotted as a function of a single variable $V$, since by symmetry all node variables can be set equal. The curves for different values of $N$ demonstrate the convexity of the potential. Each curve exhibits a global minimum corresponding to the fixed point in Eq. (11), which for small $\delta$ can be approximated as $V^{*}\approx\sqrt{K}$. b) Cost function around the fixed point $V^{*}$ for the ZKC network [29], plotted as a function of the fitness value $V_{i}$ for specific nodes. Each curve shows how the potential varies around the fixed point $V_{i}^{*}$ for that node. Nodes with higher fitness display larger values of $V_{i}^{*}$, i.e., their minima lie farther to the right. The legend reports the node degrees $d$ for comparison. As shown, the highest fitness nodes contribute most to the total potential, while low-fitness nodes provide only local corrections. Each curve exhibits a unique, global minimum. The curves are convex, though the logarithmic scale on the vertical axis may create a misleading appearance of concavity in some regions.
To establish the uniqueness of the solution to the NHEFC equations, we investigate the properties of the potential function $U(\vec{V})$ defined in Eq. (7). We consider the form of the potential $U(\vec{Z})$ expressed in terms of the natural fitness variables $Z_{i}=-\log V_{i}$, as derived in Appendix A:
@@
A fundamental result from multivariate calculus states that if the Hessian matrix of a function is positive definite throughout its domain, the function is strictly convex. Therefore, the potential function $U(\vec{Z})$ is strictly convex. A strictly convex function possesses, at most, one point where its gradient vanishes. This stationary point corresponds precisely to the condition $\dot{Z}_{i}=0$ in the dynamical system formulation (Eq. (19)), which in turn defines the fixed point of the EFC iterative map Eq. (18). Consequently, the strict convexity of $U(\vec{Z})$ guarantees that this fixed point is not only a minimum but the unique global minimum of the potential function, ensuring the uniqueness of the NHEFC solution.
The role of $\delta$ can also be analyzed through its influence on the cost function bounds. From Eq. (4), using $k_{\max}$ as the maximum node degree in the network, we obtain the constraint $\delta\leq V_{i}\leq\delta+k_{\max}/\delta$. In terms of the logarithmic variables $Z_{i}$, this translates to the bounds $-\log(\delta+k_{\max}/\delta)\leq Z_{i}\leq-\log(\delta)$. By incorporating these inequalities into Eq. (20) and retaining only the leading terms, we establish the global bounds on the potential function,
$-N\log(1/\delta)\leq U\leq E/\delta^{2},$ (10)
@@ 3.2.2 Solution in the case of two connected nodes
The specific case of a network with only two connected nodes ($N=2$) offers a tractable scenario where the algorithm’s convergence can be analyzed in full detail (refer to Appendix B for a complete treatment). By examining the dynamics over two consecutive steps, it is found that each component $V_{j}$ evolves independently according to the same one-dimensional iterative map: $V^{(n+2)}_{j}=f(V^{(n)}_{j})$, where the function is defined as $f(x)=\delta+x/(1+\delta x)$. For any positive value of the parameter $\delta$, this function $f(x)$ has derivative $10<f^{\prime}(x)<$ 1 for $x>0$, guaranteeing the existence of a unique, globally attracting fixed point within the positive domain. This fixed point is given by:
@@ 3.2.5 Reformulation of the map via gradient descent
Figure 3: Efficient convergence of the conservative map compared to the original map. a) Trajectories of the potential $U(\vec{V})$ (Eq. (7)) for a complete graph with $N=10$. Following the original map, the dynamics exhibit strong period-2 oscillations, indicating algorithm instability. In contrast, the trajectory computed by descending the cost function gradient using the conservative map (Eq. (12)) shows no oscillatory behavior and converges more efficiently to the fixed point. It requires significantly fewer steps to reach a stable solution. b) Trajectories of the potential for the ZKC network. Also in this case, the conservative map converges considerably faster and avoids oscillations. These results suggest that, for practical computations, the original map defined by Eq. (4) should be replaced with the new formulation in Eq. (12).
@@ 4 Discussion
This study establishes the mathematical foundations of two prominent economic complexity measures, ECI and NHEFC, by reformulating them as optimization problems. This approach unifies these algorithms with well-established network science techniques and provides a deeper theoretical understanding and significant practical improvements. Our findings demonstrate that the iterative algorithms for both ECI and NHEFC can be derived as procedures to minimize specific cost functions, revealing the underlying mechanics of each framework.
@@
A key result of our analysis is the contrasting nature of the cost functions associated with ECI and NHEFC. The ECI cost function is a quadratic form related to the network’s Laplacian matrix, analogous to a classical harmonic potential. This formulation promotes attraction, where connected nodes tend to have similar values, which aligns with the original ECI’s established interpretation as a spectral algorithm for network partitioning. In stark contrast, the NHEFC cost function features interaction terms involving the inverse of the node variables, resulting in a locally repulsive dynamic. This repulsive mechanism promotes differentiation among neighboring nodes. This explains the effectiveness of fitness centrality in identifying structural vulnerabilities, as it tends to assign high fitness values to nodes connected to weaker ones. More broadly, it provides a rigorous energetic perspective on the fitness solution, as initially suggested in previous work.
@@
Furthermore, our derivation of the NHEFC cost function clarifies the roles of its components. The function reveals that the natural scale of the fitness variables is logarithmic, a property observed empirically but not formally derived. Crucially, the analysis highlights the role of the regularization parameter $\delta$. This parameter acts as a damping mechanism that ensures the algorithm’s convergence by preventing the divergence caused by the hyperbolic nature of the interactions. Theoretically, the most significant contribution is the proof of the uniqueness of the NHEFC solution. By transforming the potential into its natural logarithmic variables and analyzing its Hessian matrix, we demonstrated that the potential is strictly convex. This strict convexity guarantees a unique global minimum, which corresponds to the algorithm’s fixed point, thus resolving any ambiguity about the uniqueness of the solution.
@@
The optimization framework also offers significant practical and methodological advancements. For NHEFC, we have shown that the dynamics expressed in logarithmic variables are irrotational. This implies that the most efficient computational path to the solution is a direct gradient descent of the potential function. This leads to a new “conservative map” (Eq. (12) in the original text) that converges significantly faster and avoids the period-2 oscillations characteristic of the original map. For practical applications, we recommend that the fitness algorithm be computed using this new formulation.
@@
Finally, this work broadens the applicability and the theoretical interpretation of economic complexity measures. By reformulating the underlying algorithms as optimization problems, we establish formal connections to broader scientific frameworks, including Kantorovich potentials in optimal transport theory and spectral methods in graph theory through the Laplacian matrix. Crucially, expressing the dynamics in a mono-partite setting removes the constraint of bipartite trade networks, allowing these measures to be applied to arbitrary graph structures. This generalization paves the way for their adoption in diverse contexts, such as ecological systems, infrastructure networks, and spatial economic analysis, significantly extending their scope and relevance.

### v2

@@ Abstract
Economic complexity algorithms aim to uncover the hidden capabilities that drive economic systems. Here, we present a fundamental reinterpretation of two of these algorithms, the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC), by reformulating them as optimization problems that minimize specific cost functions. We show that ECI computation is equivalent to finding eigenvectors of the network’s transition matrix by minimizing the quadratic form associated with the network’s Laplacian, thus revealing its limitations in capturing the diversification of countries. For EFC, we derive a novel cost function that exploits the algorithm’s intrinsic logarithmic structure and clarifies the role of its regularization parameter. Additionally, we establish the uniqueness of its solution, providing theoretical foundations for its application. This optimization–based reformulation bridges economic complexity and established frameworks in spectral theory, network science, and optimization. The theoretical insights translate into practical computational advantages: we introduce a conservative, gradient-based update rule that substantially accelerates algorithmic convergence, with potential implications for a broader class of algorithms, including the Sinkhorn–Knopp method. Beyond advancing our theoretical understanding of economic complexity indicators, this work opens new pathways for algorithmic improvements and extends applicability to general network structures beyond traditional bipartite economic networks.
@@ 1 Introduction
Economic complexity measures are designed to characterize the productive capabilities of countries and the structure of international trade [1, 2, 3]. A first attempt to formalize this idea was made by Hidalgo and Hausmann through the Economic Complexity Index (ECI) [4, 5], based on a linear spectral method applied to the country-product network [6, 7]. The algorithm was later shown to be mathematically equivalent to a spectral clustering method that partitions countries and products into groups. Subsequently, Tacchella et al. [1, 8] proposed a different approach with the Economic Fitness and Complexity (EFC) method, introducing a non-linear iterative framework to better capture the interplay between country diversification and product competitiveness [9, 10]. The concept of fitness is closely linked to the notion of nestedness in bipartite networks [11], and has been shown to be a reliable predictor of GDP growth [10]. Later refinements of EFC, such as the regularized Non-Homogeneous Economic Fitness and Complexity (NHEFC) method [12], aimed to ensure robust convergence and improve its applicability to empirical data.
@@
While the Economic Complexity approach has demonstrated strong empirical relevance in understanding economic growth and trade networks, its mathematical foundations remain an active area of research. Previous studies [6, 7] have explored the connections between economic complexity indices and spectral methods in network science. In particular, [13] recasts both ECI and EFC within a common spectral framework based on bipartite networks, highlighting their structural similarities despite substantial differences in algorithmic formulation. Indeed, the two approaches produce markedly different country rankings, underscoring the distinct informational content captured by each method [9]. In [14, 15], the authors investigate the relation between economic complexity indices and standard notions of centrality, showing that these algorithms can provide valuable insight into the spectral properties of the underlying graph. More recently, ECI and EFC have also been shown to be formally equivalent to network centrality measures [16]. Extending these algorithms to arbitrary network structures, beyond the canonical bipartite formulation, has significantly broadened their applicability and interpretability. Therefore, a deeper theoretical understanding of these approaches is crucial to expanding their potential applications, as already demonstrated in fields such as ecology [17, 18, 19].
@@
Building upon these developments, recent studies have proposed a connection between the EFC algorithm and optimal transport theory [20], particularly in relation to the Sinkhorn–Knopp algorithm [21], which is used to compute doubly stochastic matrices. In this context, economic complexity indices can be interpreted as solutions to optimization problems analogous to those found in optimal transport [22, 23], where cost functions play a central role in defining the equilibrium distributions [24]. Notably, the iterative structure of the EFC algorithm resembles the scaling procedures used in transport problems, and its associated cost function is associated with Kantorovich potentials [25], which characterize the optimal allocation of resources over networks. More generally, the interpretation of economic complexity algorithms as optimization problems [26, 27, 28] provides a unified framework that clarifies their mathematical structure, reveals properties of their solutions, and facilitates connections with other domains.
@@
This study demonstrates that both ECI and EFC can be reformulated as optimization problems, each minimizing a distinct cost function. Specifically, we show that the ECI algorithm minimizes an energy function derived from the network’s Laplacian matrix. In contrast, we prove that the cost function underlying EFC is strictly convex, thus providing a rigorous foundation for its optimization-based formulation. We also establish convergence to the fixed point (corresponding to the minimum of the cost function) in the case of regular graphs, and prove local convergence for general undirected and unweighted graphs. In addition, we propose a modification to the iterative algorithm that directly follows the gradient descent of the potential, significantly accelerating numerical convergence and potentially applicable to the Sinkhorn–Knopp class of algorithms for matrix renormalization.
@@ 2.2 Previous work
In our previous work [16, 24], we described two methods for quantifying economic complexity: the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC). Both approaches are based on the country–product binary matrix $\mathbf{M}$. In this matrix, the entry $M_{cp}$ is set to 1 if country $c$ significantly exports product $p$, and 0 otherwise, as determined by the Revealed Comparative Advantage (RCA) criterion [31].
@@
In contrast, EFC employs a non-linear map [1]. To ensure robust convergence, we used the regularized Non-Homogeneous version of EFC (NHEFC) given by [12]
@@
By introducing the adjacency matrix $\mathbf{A}$ and the diagonal matrix $\mathbf{K}$ with $k_{c}$ and $k_{p}$ on its diagonal, we can provide a mono-partite representation that recasts both ECI and EFC as iterative maps over a unified vector $\vec{V}$ of dimension $N=N_{c}+N_{p}$.
@@
thereby generalizing the approach, originally designed for bipartite structures, to any simple-graph structure. In Eq. (3), the notation $(\vec{V})^{-1}$ stands for taking the reciprocal of each element of the vector $\vec{V}$. This mono-partite formulation is a key generalization, allowing the ECI and EFC concepts to be applied to any arbitrary network structure, not just the bipartite country-product networks for which they were originally designed.
@@ 2.3 Cost Functions
In this section, we show that both ECI and EFC can be understood as optimization algorithms, as they can be reformulated as problems of minimizing specific cost functions. In many optimization settings, the cost (or objective) function is defined from the outset, and iterative procedures are devised to find its minimum. In our case, the process is reversed: starting from the iterative maps in Eq. (2) and Eq. (3), we reconstruct the corresponding cost functions by integration. The cost function associated with ECI relates directly to the standard eigenvector problem and can be interpreted within that framework. For EFC, the connection to optimization has been previously noted due to structural similarities between the fitness algorithm and the Sinkhorn–Knopp algorithm from optimal transport theory [32]. However, while prior work has identified this formal analogy [20], our study provides the first explicit derivation of EFC’s cost function. This explicit formulation enables us to prove the function’s convexity rigorously and subsequently develop a more efficient gradient-based algorithm. In the following, we derive the explicit expression of the cost function associated with the EFC algorithm. Full details of the derivations for both ECI and EFC are provided in Appendix A.
@@ 3 Results
In this section, we analyze the convergence properties of ECI and EFC. For ECI, properties such as the convexity of the potential and the existence of a fixed point follow directly from the fact that ECI is equivalent to computing the second eigenvector of the matrix $\mathbf{K}^{-1}\mathbf{A}$. The following part focuses on the convergence of EFC across different network structures, starting from a simple two-node undirected graph and extending to regular graphs and arbitrary topologies. The analysis establishes local convergence criteria near fixed points and explores the influence of the non-homogeneous parameter $\delta$ on the algorithm’s behavior. Finally, we introduce an alternative formulation of the fitness algorithm that follows the gradient of the potential function. This conservative formulation defines the most direct path to the fixed point and significantly reduces the number of iterations required for convergence.
@@ 3.1 Convergence properties of ECI
Since this fixed point (the constant vector) does not provide a meaningful solution of the problem of ranking countries or products, the original formulation of ECI [4] prescribes stopping the iteration after a fixed number of steps. As discussed in several works [13, 24, 2], the intermediate steps of the iteration retain information about the subdominant eigenvectors, particularly the second one. As a result, ECI can be obtained in two equivalent ways: (i) by iterating the map Eq. (2) for a fixed number of steps, or (ii) by directly computing the second eigenvector of the transition matrix $\mathbf{K}^{-1}\mathbf{A}$. This eigenvector maximizes the partitioning of the network, in line with the interpretation of ECI as a spectral algorithm for community detection [13]. This observation further supports the evidence that ECI has intrinsic limitations in capturing the true diversification structure of countries [9].
3.2 Convergence properties of EFC
@@ 3.2.1 Uniqueness of solution via convexity analysis
Figure 1: Energetic landscape of the fitness algorithm. a) Cost function from Eq. (7) for regular complete graphs with varying number of nodes $N$ and K=N1K=N-1. The potential $U(V)$ is plotted as a function of a single variable $V$, since by symmetry all node variables can be set equal. The curves for different values of NN demonstrate the convexity of the potential. Each curve exhibits a global minimum corresponding to the fixed point in Eq. (11), which for small $\delta$ can be approximated as $V^{*}\approx\sqrt{K}$. b) Cost function around the fixed point $V^{*}$ for the ZKC network [30], plotted as a function of the fitness value $V_{i}$ for specific nodes. Each curve shows how the potential varies around the fixed point $V_{i}^{*}$ for that node. Nodes with higher fitness display larger values of $V_{i}^{*}$, i.e., their minima lie farther to the right. The legend reports the node degrees $d$ for comparison. As shown, the highest fitness nodes contribute most to the total potential, while low-fitness nodes provide only local corrections. Each curve exhibits a unique, global minimum. The curves are convex, though the logarithmic scale on the vertical axis may create a misleading appearance of concavity in some regions.
To establish the uniqueness of the solution to the EFC equations, we investigate the properties of the potential function $U(\vec{V})$ defined in Eq. (7). We consider the form of the potential $U(\vec{Z})$ expressed in terms of the natural fitness variables $Z_{i}=-\log V_{i}$, as derived in Appendix A:
@@
A fundamental result from multivariate calculus states that if the Hessian matrix of a function is positive definite throughout its domain, the function is strictly convex. Therefore, the potential function $U(\vec{Z})$ is strictly convex. A strictly convex function possesses, at most, one point where its gradient vanishes. This stationary point corresponds precisely to the condition $\dot{Z}_{i}=0$ in the dynamical system formulation (Eq. (19)), which in turn defines the fixed point of the EFC iterative map Eq. (18). Consequently, the strict convexity of $U(\vec{Z})$ guarantees that this fixed point is not only a minimum but the unique global minimum of the potential function, ensuring the uniqueness of the EFC solution.
The role of $\delta$ can also be analyzed through its influence on the cost function bounds. From Eq. (4), using $k_{\max}$ as the maximum node degree in the network, we obtain the constraint $\delta\leq V_{i}\leq\delta+k_{\max}/\delta$. In terms of the logarithmic variables $Z_{i}$, this translates to the bounds $\log(\delta+k_{\max}/\delta)\leq Z_{i}\leq-\log(\delta)$. By incorporating these inequalities into Eq. (20) and retaining only the leading terms, we establish the global bounds on the potential function,
$N\log(1/\delta)\leq U\leq E/\delta^{2},$ (10)
@@ 3.2.2 Solution in the case of two connected nodes
The specific case of a network with only two connected nodes ($N=2$) offers a tractable scenario where the algorithm’s convergence can be analyzed in full detail (refer to Appendix B for a complete treatment). By examining the dynamics over two consecutive steps, it is found that each component $V_{j}$ evolves independently according to the same one-dimensional iterative map: $V^{(n+2)}_{j}=f(V^{(n)}_{j})$, where the function is defined as $f(x)=\delta+x/(1+\delta x)$. For any positive value of the parameter $\delta$, this function $f(x)$ has derivative $0<f^{\prime}(x)<1$ for $x>0$, guaranteeing the existence of a unique, globally attracting fixed point within the positive domain. This fixed point is given by:
@@ 3.2.5 Reformulation of the map via gradient descent
Figure 3: Efficient convergence of the conservative map compared to the original map. a) Trajectories of the potential $U(\vec{V})$ (Eq. (7)) for a complete graph with $N=10$. Following the original map, the dynamics exhibit strong period-2 oscillations, indicating algorithm instability. In contrast, the trajectory computed by descending the cost function gradient using the conservative map (Eq. (12)) shows no oscillatory behavior and converges more efficiently to the fixed point. It requires significantly fewer steps to reach a stable solution. b) Trajectories of the potential for the ZKC network. Also in this case, the conservative map converges considerably faster and avoids oscillations. Due to the formal analogy between EFC and the Sinkhorn–Knopp algorithm [21], the conservative map offers a more efficient alternative for matrix renormalization.
@@
Given the formal equivalence between the EFC algorithm and the Sinkhorn–Knopp algorithm [20, 21], which is used for matrix renormalization and the computation of doubly stochastic matrices, the conservative map we propose can also be seen as an efficient alternative in this broader algorithmic class.
@@ 4 Discussion
This study establishes the mathematical foundations of two economic complexity measures, ECI and EFC, by reformulating them as optimization problems. This approach unifies these algorithms with well-established network science techniques and provides a deeper theoretical understanding and significant practical improvements. Our findings demonstrate that the iterative algorithms for both ECI and EFC can be derived as procedures to minimize specific cost functions, revealing the underlying mechanics of each framework.
@@
A key result of our analysis is the contrasting nature of the cost functions associated with ECI and EFC. The ECI cost function is a quadratic form related to the network’s Laplacian matrix, analogous to a classical harmonic potential. This formulation promotes attraction, where connected nodes tend to have similar values, which aligns with the established interpretation of ECI as a spectral algorithm for network partitioning. In stark contrast, the EFC cost function features interaction terms involving the inverse of the node variables, resulting in a locally repulsive dynamic. This repulsive mechanism promotes differentiation among neighboring nodes. This explains the effectiveness of fitness centrality in identifying structural vulnerabilities, as it tends to assign high fitness values to nodes connected to weaker ones. More broadly, it establishes a rigorous connection with well-known optimization problems and contributes to understanding the relationship between Economic Fitness and Complexity and the broader framework of optimal transport theory.
@@
Furthermore, our derivation of the EFC cost function clarifies the roles of its components. The function reveals that the natural scale of the fitness variables is logarithmic, a property observed empirically but not formally derived. Crucially, the analysis highlights the role of the regularization parameter $\delta$. This parameter acts as a damping mechanism that ensures the algorithm’s convergence by preventing the divergence caused by the hyperbolic nature of the interactions. Theoretically, the most significant contribution is the proof of the uniqueness of the EFC solution. By transforming the potential into its natural logarithmic variables and analyzing its Hessian matrix, we demonstrated that the potential is strictly convex. This strict convexity guarantees a unique global minimum, which corresponds to the algorithm’s fixed point, thus resolving any ambiguity about the uniqueness of the solution.
@@
The optimization framework also offers significant practical and methodological advancements. For EFC, we have shown that the dynamics expressed in logarithmic variables are irrotational. This implies that the most efficient computational path to the solution is a direct gradient descent of the potential function. This leads to a new “conservative map” (Eq. (12) in the original text) that converges significantly faster and avoids the period-2 oscillations characteristic of the original map. Moreover, given the formal connection between the EFC algorithm and the Sinkhorn–Knopp algorithm, used for matrix renormalization and the computation of doubly stochastic matrices, the conservative map introduced here may also provide an effective computational strategy within that broader class of problems. Future work should explore whether this gradient-based formulation can systematically improve the convergence speed, stability, and theoretical understanding of Sinkhorn–type algorithms, particularly in the context of optimal transport and large-scale network renormalization.
@@
Finally, this work broadens the applicability and the theoretical interpretation of economic complexity measures. By reformulating the underlying algorithms as optimization problems, we establish formal connections to broader scientific frameworks, including Kantorovich potentials in optimal transport theory and spectral methods in graph theory through the Laplacian matrix. Crucially, expressing the dynamics in a mono-partite setting removes the constraint of bipartite trade networks, allowing these measures to be applied to arbitrary graph structures. This generalization opens the way for their adoption in diverse contexts, such as ecological systems, infrastructure networks, and spatial economic analysis, significantly extending their scope and relevance.